OpenStudy (anonymous):

Find the length of the hypotenuse of an isosceles right triangle whose legs are 1 unit in length.

OpenStudy (welshfella):

this must be an isosceles trianggle which is also right angled

OpenStudy (welshfella):



OpenStudy (welshfella):

h^2 = 1^2 + 1^2

OpenStudy (welshfella):

h^2 = 1 + 1